Ten pounds of lemon drops are to be mixed with cherry chews to make a mixture that will sell for $1.80 per pound. How much of the more expensive candy should be used? See the table
Candy Price per pound
Peppermint patties $1.35
Lemon drops $1.60
Licorice lumps $1.75
Cherry chews $1.90

Answer by richwmiller(17219)

1.9*x+1.6*10=1.8(10+x)
1.9*x+16=18+1.8x
1.9x-1.8x=18-16
0.1x=2
x=20 is amount at chews and 30 is amount of mix
check
1.9*20+1.6*10=1.8(10+20)
38+16=1.8(30)
54=54
ok